# On the generalization of language models from in-context learning and finetuning: a controlled study

### v1

@@ Discussion
In this paper, we have performed controlled experiments on how language models generalize various types of novel information from in-context learning and finetuning. Overall, we find that the models generalize better on average along several dimensions from in-context learning. Using in-context learning to augment the finetuning dataset can exploit the complementary benefits of both to yield better performance.
The distinct inductive biases of in-context learning and finetuning: A variety of works have studied the inductive biases of in-context learning. One common theme has been to emphasize that in-context learning can approximate gradient descent (e.g. Von Oswald et al., 2023), in settings where that behavior is optimal. However, a variety of other works have found that the inductive biases of in-context learning can vary depending on factors such as dataset diversity (Raventós et al., 2024) or model scale (Wei et al., 2023). Several works have explicitly noted the distinct inductive biases of in-context learning and finetuning (Chan et al., 2022b; Shen et al., 2023). Our work contributes to this line of findings.
Accessible information and learning by thinking: Lombrozo (2024) highlights how “learning by thinking” is a unifying theme across cognitive science and recent advances in AI—a system can improve performance through computation without further inputs. Lombrozo highlights that while superficially this may seem paradoxical—information cannot be created—this further computation can increase the accessibility of information and thus improve performance in practice. This argument parallels the theoretical one made by Xu et al. (2020) on how computation can increase the accessibility of information. Our use of in-context inferences to improve finetuning performance beyond the original data follows this pattern. For example, the information about reversals and syllogisms is always hidden within the data, but finetuning on the in-context inferences make this information more explicit and thus more readily accessible at test time.
@@
Conclusions: We have explored the generalization of in-context learning and finetuning when LMs are exposed to various types of novel information structures. We find notable differences—often with ICL generalizing better to certain types of inferences—and propose ways of achieving better performance by adding in-context inferences to finetuning data. We hope this work will contribute to the science of understanding learning and generalization in foundation models, and the practicalities of adapting them to downstream tasks.

### v2

@@ Discussion
In this paper, we performed controlled experiments on how language models generalize various types of novel information from in-context learning and finetuning. Overall, we find that the models generalize better on average along several dimensions from in-context learning. Using in-context learning to augment the finetuning dataset can exploit the complementary benefits of both to yield better performance.
The distinct inductive biases of in-context learning and finetuning: A variety of works have studied the inductive biases of in-context learning. One common theme has been to emphasize that in-context learning can approximate gradient descent (e.g. Von Oswald et al., 2023), in settings where that behavior is optimal. However, a variety of other works have found that the inductive biases of in-context learning can vary depending on factors such as dataset diversity (Raventós et al., 2024) or model scale (Wei et al., 2023). Indeed, Elmoznino et al. (2024) argue that in-context learning Several works have explicitly noted the distinct inductive biases of in-context vs. in-weights learning (e.g. “rule-based” vs “exemplar-based” generalization) (Chan et al., 2022b; Shen et al., 2023; Awadalla et al., 2022)—while others have argued that they are more similar than they might seem (Mosbach et al., 2023). Our work contributes to this line of findings, but goes beyond the typical input-output tasks considered in these works, to consider other types of knowledge.
Accessible information and learning by thinking: Lombrozo (2024) highlights how “learning by thinking” is a unifying theme across cognitive science and recent advances in AI—a system can acquire new information and skills purely through computation, without further inputs. Lombrozo highlights that while superficially this may seem paradoxical—information cannot be created—this further computation can increase the accessibility of information and thus improve performance in practice. This argument parallels the theoretical one made by Xu et al. (2020) on how computation can increase the accessibility of information. Our use of in-context inferences to improve finetuning performance beyond the original data follows this pattern. For example, the information about reversals and syllogisms is always hidden within the data, but finetuning on the in-context inferences make this information more explicit and thus more readily accessible at test time.
@@
Conclusions: We have explored the generalization of in-context learning and finetuning when LMs are exposed to various types of novel information structures. We find notable differences—often with ICL generalizing better for certain types of inferences—and propose ways of achieving better performance by adding in-context inferences to finetuning data. We hope this work will contribute to the science of understanding learning and generalization in foundation models, and the practicalities of adapting them to downstream tasks.